# CityLens: Benchmarking Large Language-Vision Models for Urban Socioeconomic Sensing

### v1

@@ Abstract
Understanding urban socioeconomic conditions through visual data is a challenging yet essential task for sustainable urban development and policy planning. In this work, we introduce CityLens, a comprehensive benchmark designed to evaluate the capabilities of large language-vision models (LLVMs) in predicting socioeconomic indicators from satellite and street view imagery. We construct a multi-modal dataset covering a total of 17 globally distributed cities, spanning 6 key domains: economy, education, crime, transport, health, and environment, reflecting the multifaceted nature of urban life. Based on this dataset, we define 11 prediction tasks and utilize three evaluation paradigms: Direct Metric Prediction, Normalized Metric Estimation, and Feature-Based Regression. We benchmark 17 state-of-the-art LLVMs across these tasks. Our results reveal that while LLVMs demonstrate promising perceptual and reasoning capabilities, they still exhibit limitations in predicting urban socioeconomic indicators. CityLens provides a unified framework for diagnosing these limitations and guiding future efforts in using LLVMs to understand and predict urban socioeconomic patterns. Our codes and datasets are open-sourced via https://github.com/tsinghua-fib-lab/CityLens.
@@ 1 Introduction
Understanding the socioeconomic characteristics of urban regions is fundamental to the planning, management, and sustainability of cities. Urban socioeconomic sensing, the process of quantifying indicators such as income, education, health, and transport conditions across spatial units, plays a critical role in shaping how cities function and evolve. These indicators directly influence residents’ quality of life and are deeply intertwined with key aspects of urban inequality, mobility, and resource allocation. Moreover, urban socioeconomic data serves as a cornerstone for measuring progress toward several United Nations Sustainable Development Goals (UN SDGs). Accurate and timely information on urban disparities is essential for tracking these goals and designing effective interventions. In practice, governments and urban planners rely on socioeconomic indicators to inform a wide range of decisions—from zoning regulations and infrastructure investment to public health strategies. A better understanding of spatially resolved urban indicators empowers decision-makers to allocate resources more equitably, respond to local needs, and promote inclusive urban development.
@@
A growing body of work has explored the use of classical deep learning methods to predict urban socioeconomic indicators. Some approaches, such as lbkgzhou , leverage knowledge graphs to infer socioeconomic indicators. In parallel, researchers have explored the use of urban imagery to understand cities through their visual appearance. Methods such as li2022predicting liu2023knowcl linurbanvillage , and yong2024musecl employ contrastive learning to generate visual representations from street view or satellite images, while others like fan2023uvi apply basic computer vision models to extract visual features. However, the classical methods face several key limitations, including difficulty in handling unstructured or multi-modal data, the inability to work across multiple countries, and cannot interpret subjective and culturally significant aspects of place. Nevertheless, Large Language-Vision Models are inherently equipped to address these challenges with their ability to integrate multiple modalities, generalize globally, and interpret cultural nuances.
@@
In recent years, researchers have begun to leverage LLVMs and large language models (LLMs) to address some of the limitations of classical approaches. For example, yan2024urbanclip and hao2025urbanvlp employ LLVM to generate textual descriptions from urban imagery, effectively introducing a textual modality to enrich visual understanding. Other studies, such as manvi2024geollm and manvi2024large , explore the ability of LLMs to predict socioeconomic indicators directly through textual prompts, and further examine issues like geographical bias across different countries. Despite these promising advances, existing works still fall short in several key aspects. Most efforts are limited in terms of spatial coverage, indicator diversity, and multi-modal integration. Crucially, there remains a lack of a systematic and unified benchmark to comprehensively evaluate how LLVMs perform across tasks, regions, and modalities in the context of urban socioeconomic sensing.
@@
To address these limitations, we propose CityLens, a comprehensive benchmark designed to evaluate the ability of large vision-language models to predict urban socioeconomic indicators using both street view and satellite imagery. CityLens spans a total of 17 cities across multiple continents, covering 11 indicators across 6 socioeconomic domains, including economy, health, education, environment, transport, and crime. By integrating diverse data modalities and global geographic coverage, CityLens enables systematic, cross-task, and cross-region evaluation of LLVMs’ capabilities in urban perception, geo-visual reasoning, and numerical estimation. Overall, our contributions are summarized as follows:
@@
  • We introduce CityLens, a multi-modal benchmark for urban socioeconomic sensing. It is built through a scalable data collection and processing pipeline, covering 17 cities across different continents, 11 indicators in 6 socioeconomic domains, using both street view and satellite imagery.
@@
  • We provide detailed empirical analyses that offer new insights into how different prompting strategies, input modalities, and task types affect model performance, highlighting challenges, opportunities, and future directions for socioeconomic sensing with LLVMs.
@@ 2 Methods
In this paper, we present CityLens, a comprehensive benchmark designed to evaluate the capabilities of large language-vision models in predicting socioeconomic indicators from both satellite and street view imagery. As illustrated in Figure 1, CityLens spans 11 real-world indicators across 6 socioeconomic domains, covering 17 globally distributed cities with diverse urban forms and development levels. To systematically assess model performance, we evaluate 17 different LLVMs using 3 distinct evaluation paradigms.
Figure 1: Framework of CityLens.
@@ 2.1 Dataset Construction
To support the evaluation of LLVMs across diverse socioeconomic indicators, as Figure 2 illustrates, we construct a region-level dataset for indicators. Each region is represented by one satellite image and street view images, which are jointly provided as input to the LLVMs.
2.1.1 Tasks and Sources
@@
We provide a detailed list of data sources for all indicators in the Appendix A.1; here, we briefly describe the collected indicators. Under the economy domain, we cover 7 critical indicators: Gross Domestic Product (GDP), house price, population, median household income, poverty 100%, poverty 200%, and income Gini coefficient. In the transport domain, we include seven indicators: PMT, VMT, PTRP, VTRP, walk and bike ratio, public transport ratio, and drive ratio. In the crime domain, we focus on two indicator: violent crime incidence and non-violent crime incidence, both defined as the number of crime occurrences per census tract. For the health domain, we include 9 kinds of indicators to capture different dimensions of urban health outcomes: obesity, diabetes, cancer, no leisre-time physical activity(LPA), mental health, physical health, depression rate, accessibility to healthcare, and life expectancy. In the environment domain, we consider two indicators: carbon emissions and building height. Building height is increasingly used as an explicit yet indirect indicator of urban socioeconomic development, population density, and land use intensity. Under the education domain, following liu2023knowcl , we use the bachelor ratio, defined as the proportion of residents holding a bachelor’s degree or higher, as the target variable. These domains are selected to ensure a balanced and holistic representation of urban conditions that are commonly studied in social science and urban planning.
2.1.2 Data Collection
@@
Since many ground-truth indicators are only available for specific countries (the US and the UK), we focus on region-level prediction tasks in three representative cities from each country. We choose New York, San Francisco, and Chicago in the US, and Leeds, Liverpool, and Birmingham in the UK. For globally available indicators, we expand coverage to cities across 6 continents, including: Cape Town, Nairobi, London, Paris, Beijing, Shanghai, Moscow, Mumbai, Tokyo, Sao Paulo, and Sydney, which ensures cross-regional evaluation diversity. Beyond ground-truth indicator data, we collect both satellite images and street view images for each task region. We obtain street view images for Beijing and Shanghai using the Baidu Maps API, while for other cities, we utilize the Google Street View API. Additionally, satellite images are downloaded from Esri World Imagery, providing zoom level 15 images for all cities in our study.
Figure 2: Benchmark Construction Pipeline, including data collection, indicator selection, task construction and evaluation method.
2.1.3 Indicator Selection
@@
2.1.4 Task Construction
For US-only indicators, we define census tract level prediction tasks. Each census tract is represented by one satellite image and 10 street view images, which are jointly input into the LLVM for prediction. While prior work such as fan2023uvi uses 20 or more street view images to represent each region, we found that this setup is often impractical for LLVMs. Specifically, we initially experiment with 20 street view images per region, but observe that this would significantly increase the computational cost, and exceed the input limits of models like Gemini, which can only process up to 10 images per inference, and also frequently hit the token length limit of other models. Therefore, we adopt a compromise of 10 images per region to ensure compatibility across models while maintaining sufficient visual context. For UK-only indicators, tasks are constructed at the MSOA level using a similar strategy. For global tasks, we first download multiple satellite images covering each city’s spatial extent. Then, we randomly sample 30 geographic points within the satellite image coverage, downloading at least 10 street view images for those points. Each satellite image is paired with its local indicator value to form an evaluation unit. For the China house price task, we follow the global task methodology and select Beijing and Shanghai as target cities. Due to resource constraints, we randomly sample up to 500 cases per task for country-specific indicators, and up to 1000 cases per task for global indicators. In practice, some tasks contain fewer samples due to data availability limitations, but these values represent the maximum sample size allowed per task. The detailed statistics of available data before applying the sampling strategy are presented in Figure 3(b).
@@ 2.1.4 Task Construction
Figure 3: (a) 11 indicators in CityLens benchmark and their case counts. (b) Statistics of the CityLens dataset, including urban images.
@@ 2.2 Evaluation Methodologies
We design three distinct paradigms to explore the capability of LLVM for socioeconomic indicator prediction. Each paradigm is aimed at assessing different facets of how LLVM can be applied to socioeconomic indicator sensing.
2.2.1 Direct Metric Prediction
Direct Metric Prediction refers to providing region-level urban imagery and directly querying the LLVM for the metric value, such as: "What’s the percentage of the population commuting by public transit in this census tract?" In addition, the prompt positions the model as an urban socioeconomic scientist in a specific city. Despite this, the model faces significant challenges in accurately predicting the exact true values of these indicators.
2.2.2 Normalized Metric Estimation
Given the difficulty of directly predicting precise indicator values, we adopt a Normalized Metric Estimation approach inspired by GeoLLM manvi2024geollm . Specifically, we transform all indicator values into a normalized range from 0.0 to 9.9, discretized to one decimal place. The model is then prompted to estimate this normalized value based on the input images. This formulation aims to investigate whether the LLVM possesses coarse-grained spatial knowledge and the ability to associate visual cues with relative indicator levels.
2.2.3 Feature-Based Regression
In the Feature-Based Regression approach, we first design a structured prompt that guides the LLVM to evaluate each street view image along 13 predefined visual attributes, following the visual taxonomy proposed by fan2023uvi . These features capture key elements of the urban environment, such as greenery, vehicle, facade, and sidewalk. For each region, we represent its visual environment using 10 sampled street view images. For each visual feature, we compute the average score across these images, resulting in a single feature vector that characterizes the region. These aggregated visual features are then used as inputs to a LASSO regression model, which is trained to predict the corresponding ground-truth indicator values using a 5-fold cross-validation setup.
@@ 3 Experiments
3.1 Settings
3.1.1 LLVMs
We consider a diverse set of LLVMs as baselines to benchmark our proposed methods. The selected models include both open-source and proprietary systems, covering a range of model sizes and capabilities. We choose Gemma3-4B/12B/27B gemma_2025 , Qwen2.5VL-3B/7B/32B Qwen2.5-VL , Llama4-Scout/Maverick llama4 , Mistral-small-3.1-24B mistral , Phi-4-multimodal abouelenin2025phi , MiniMax-01 li2025minimax , Gemini-2.0-flash/Gemini-2.5-flash gemini2.0 , GPT-4o-mini achiam2023gpt , GPT-4.1-mini/nano gpt4.1 and Amazon-Nova-Lite amazonnova . One thing to note is that models in the Gemini series can accept at most 10 images as input. Therefore, for this series, we use 1 satellite image and 9 street view images per region to stay within the model’s input constraints.
3.1.2 Metrics
@@
Tasks GDP Pop. HP VC PT DR BH MH AH LF BR Mean SD
@@
3.2 Overall Performance on Feature-Based Regression
3.2.1 Challenge of the Benchmark for LLVMs
As shown in Table 1, the overall performance suggests that our benchmark poses a significant challenge for current large language-vision models. In particular, tasks such as Mental Health and Bachelor Ratio exhibit low $R^{2}$ scores, in some cases even approaching zero, e.g., 0.001. This highlights the difficulty of CityLens in Feature-Based Regression method: even when leveraging visual features extracted by advanced LLVMs, the resulting representations often fail to capture the complex patterns required for accurate prediction of socioeconomic indicators, especially in domains like health and education, where visual cues are often subtle, noisy, or indirect.
3.2.2 Performance Differences Across Models
We observe substantial performance differences across LLVMs, reflecting how model scale, architecture, and training design influence their ability to extract meaningful visual features for downstream prediction. Comparing models within the same series but at different scales, we find that increasing model size does not always guarantee better performance. For example, Gemma3-12B achieves the best score on GDP and Life Expectancy, yet the 27B variant, despite its larger size, performs worse in these two tasks, with relative drops of 4.3% and 6.8% respectively. This counterintuitive result may be attributed to the unique nature of socioeconomic sensing tasks, which requires the model to consistently extract and score a predefined set of nuanced visual features from urban imagery. When comparing models from different series with similar parameter scales, clear differences emerge. For instance, Gemma3-4B significantly outperforms Qwen2.5VL-3B in nearly all tasks, with relative improvements ranging from 6.4% to 255% across different indicators, suggesting that Gemma’s architecture or training process may enable more consistent and informative scoring of urban visual features, which in turn leads to better performance in downstream socioeconomic prediction.
3.2.3 Variations Across Different Task Types
Performance also varies across task types. Tasks like Building Height, Public Transport, and GDP tend to have relatively higher values across models, with Building Height reaching an $R^{2}$ of 0.59, suggesting that these indicators are associated with more observable visual cues that can be directly captured from street view images. For instance, Building Height is closely linked to the skyline and the vertical structure visible in images; Public Transport usage may be inferred from the presence of bus stops, transit signs, or road markings. In contrast, tasks such as Life Expectancy and Mental Health remain highly challenging, exhibiting low or near-zero predictive scores for many models. These indicators are influenced by latent factors such as lifestyle, stress levels, or social cohesion, which do not have clear or direct visual manifestations in the urban environment. Even if certain proxies exist, such as the presence of graffiti or the amount of green space, they are often subtle or semantically ambiguous, making it difficult for current LLVMs to interpret them reliably and consistently.

### v2

# arXiv:2506.00530v2[cs.AI] 01 Mar 2026

@@ Abstract
Understanding urban socioeconomic conditions through visual data is a challenging yet essential task for sustainable urban development and policy planning. In this work, we introduce CityLens, a comprehensive benchmark designed to evaluate the capabilities of Large Vision-Language Models (LVLMs) in predicting socioeconomic indicators from satellite and street view imagery. We construct a multi-modal dataset covering a total of 17 globally distributed cities, spanning 6 key domains: economy, education, crime, transport, health, and environment, reflecting the multifaceted nature of urban life. Based on this dataset, we define 11 prediction tasks and utilize 3 evaluation paradigms: Direct Metric Prediction, Normalized Metric Estimation, and Feature-Based Regression. We benchmark 17 state-of-the-art LVLMs across these tasks. These make CityLens the most extensive socioeconomic benchmark to date in terms of geographic coverage, indicator diversity, and model scale. Our results reveal that while LVLMs demonstrate promising perceptual and reasoning capabilities, they still exhibit limitations in predicting urban socioeconomic indicators. CityLens provides a unified framework for diagnosing these limitations and guiding future efforts in using LVLMs to understand and predict urban socioeconomic patterns. The code and data are available at https://github.com/tsinghua-fib-lab/CityLens.
22footnotetext: Corresponding author, email: fengj12ee@hotmail.com, liyong07@tsinghua.edu.cn, panhui@ust.hk.
@@ 1 Introduction
Understanding the socioeconomic characteristics of urban regions is fundamental to the planning, management, and sustainability of cities. Urban socioeconomic sensing, the process of quantifying indicators such as income, education, health, and transport conditions across spatial units, plays a critical role in shaping how cities function and evolve. These indicators directly influence residents’ quality of life and are deeply intertwined with key aspects of urban inequality, mobility, and resource allocation. Moreover, urban socioeconomic data serves as a cornerstone for measuring progress toward several United Nations Sustainable Development Goals (UN SDGs) (Nations, 2015). Accurate and timely information on urban disparities is essential for tracking these goals and designing effective interventions. In practice, governments and urban planners rely on socioeconomic indicators to inform a wide range of decisions—from zoning regulations and infrastructure investment to public health strategies. A better understanding of spatially resolved urban indicators empowers decision-makers to allocate resources more equitably, respond to local needs, and promote inclusive urban development.
@@
A growing body of work has explored the use of classical deep learning methods to predict urban socioeconomic indicators. Some approaches, such as Zhou et al. (2023), leverage knowledge graphs to infer socioeconomic indicators. In parallel, researchers have explored the use of urban imagery to understand cities through their visual appearance. Methods such as Li et al. (2022)Liu et al. (2023b)Lin et al. (2024), and Yong & Zhou (2024) employ contrastive learning to generate visual representations from street view or satellite images, while others apply basic computer vision models to extract visual features (Fan et al., 2023; Zhang et al., 2024). However, the classical methods face several key limitations, including difficulty in handling unstructured or multi-modal data, the inability to work across multiple countries, and cannot interpret subjective and culturally significant aspects of place. Nevertheless, large vision-language models are inherently equipped to address these challenges with their ability to integrate multiple modalities, generalize globally, and interpret cultural nuances.
@@
In recent years, researchers have begun to leverage LVLMs and large language models (LLMs) to address some of the limitations of classical approaches (Hou et al., 2025; Yan et al., 2024; Hao et al., 2025; Manvi et al., 2024b; a). For example, Yan et al. (2024) and Hao et al. (2025) employ LVLM to generate textual descriptions from urban imagery, effectively introducing a textual modality to enrich visual understanding. Other studies, such as Manvi et al. (2024b; a), explore the ability of LLMs to predict socioeconomic indicators directly through textual prompts, and further examine issues like geographical bias across different countries. Despite these promising advances, existing works still fall short in several key aspects. Most efforts are limited in terms of spatial coverage, indicator diversity, and multi-modal integration. Crucially, there remains a lack of a systematic and unified benchmark to comprehensively evaluate how LVLMs perform across tasks, regions, and modalities in the context of urban socioeconomic sensing.
Figure 1: CityLens comprises 11 socioeconomic indicators spanning six key urban domains.
@@
To address these limitations, we propose CityLens, a comprehensive benchmark designed to evaluate the ability of large vision-language models to predict urban socioeconomic indicators using both street view and satellite imagery. CityLens spans a total of 17 cities across multiple continents, covering 11 indicators across 6 socioeconomic domains, including economy, health, education, environment, transport, and crime. By integrating diverse data modalities and global geographic coverage, CityLens enables systematic, cross-task, and cross-region evaluation of LVLMs’ capabilities in urban perception, geo-visual reasoning, and numerical estimation. Overall, our contributions are summarized as follows:
@@
  • To the best of our knowledge, CityLens is the largest benchmark in urban socioeconomic sensing, along geographic coverage, indicator diversity, and model scale. It covers 17 cities across different continents, 11 indicators in 6 socioeconomic domains, using both street view and satellite imagery.
@@
  • We design extensive experiments and provide detailed analysis that offers new insights into how input configuration, model architecture, and task design affect model performance, highlighting challenges, opportunities, and future directions for socioeconomic sensing with LVLMs.
@@ 2 Methods
In this paper, we present CityLens, a comprehensive benchmark designed to evaluate the capabilities of large vision-language models in predicting socioeconomic indicators from both satellite and street view imagery. As illustrated in Figure 1, CityLens spans 11 real-world indicators across 6 socioeconomic domains, covering 17 globally distributed cities with diverse urban forms and development levels. To systematically assess model performance, we evaluate 17 different LVLMs using 3 distinct evaluation paradigms.
@@ 2.1 Dataset Construction
To support the evaluation of LVLMs across diverse socioeconomic indicators, as Figure 2 illustrates, we construct a region-level dataset by performing data collection, indicator selection, and data mapping. Each region is represented by 1 satellite image and 10 street view images, and is associated with corresponding socioeconomic indicator values.
Data Collection
@@
We provide a detailed list of data sources for all indicators in the Appendix A.5.1; here, we briefly describe the collected indicators. Under the economy domain, we cover 7 critical indicators: Gross Domestic Product (GDP), house price, population, median household income, poverty 100%, poverty 200%, and income Gini coefficient. In the transport domain, we include seven indicators: PMT, VMT, PTRP, VTRP, walk and bike ratio, public transport ratio, and drive ratio. In the crime domain, we focus on two indicator: violent crime incidence and non-violent crime incidence, both defined as the number of crime occurrences per census tract. For the health domain, we include 9 kinds of indicators to capture different dimensions of urban health outcomes: obesity, diabetes, cancer, no leisre-time physical activity(LPA), mental health, physical health, depression rate, accessibility to healthcare, and life expectancy. In the environment domain, we consider two indicators: carbon emissions and building height. Building height is increasingly used as an explicit yet indirect indicator of urban socioeconomic development, population density, and land use intensity. Under the education domain, following Liu et al. (2023b), we use the bachelor ratio, defined as the proportion of residents holding a bachelor’s degree or higher, as the target variable. These domains are selected to ensure a balanced and holistic representation of urban conditions that are commonly studied in social science and urban planning.
@@
Since many ground-truth indicators are only available for specific countries (the US and the UK), we focus on region-level prediction tasks in three representative cities from each country. We choose New York, San Francisco, and Chicago in the US, and Leeds, Liverpool, and Birmingham in the UK. For globally available indicators, we expand coverage to cities across 6 continents, including Cape Town, Nairobi, London, Paris, Beijing, Shanghai, Moscow, Mumbai, Tokyo, Sao Paulo, and Sydney, which ensures cross-regional evaluation diversity. Beyond ground-truth indicator data, we collect both satellite images and street view images for each task region. We obtain street view images for Beijing and Shanghai using the Baidu Maps API, while for other cities, we utilize the Google Street View API. All experimental results reported in the main paper are based on these Google and Baidu sourced street view images. To promote transparency, completeness, and reproducibility of CityLens, we further construct an alternative version of the dataset using publicly accessible street view images from Mapillary, referred to as CityLens-Mapillary. We report the benchmark results based on Mapillary street view images in Appendix A.3. Additionally, the $\times256$-pixel satellite images with about 4.7 m-resolution are downloaded from Esri World Imagery.
Indicator Selection
@@
Figure 2: Benchmark Construction Pipeline, including data collection, indicator selection, data mapping and evaluation methods.
Data Mapping
In CityLens, each region serves as a prediction unit, represented by 1 satellite image and 10 street view images, and is paired with a set of scalar labels corresponding to multiple target indicators. These labels are computed by mapping and aggregating raw tabular data from heterogeneous sources to the respective region. As shown in the region column of Figure 3(b), we define census tract-level prediction tasks for US-only indicators, and construct MSOA-level tasks for UK-only indicators using a similar strategy. For global tasks, each satellite image coverage area constitutes an evaluation unit. We first download multiple satellite images covering each city’s spatial extent. Then, within each satellite image’s coverage, we randomly sample 30 geographic points and collect at least 10 street view images corresponding to those points. For the China house price task, we follow the global task methodology and select Beijing and Shanghai as target cities. A detailed explanation of the mapping and aggregating of data from various sources and geographic scales for each task is provided in Appendix A.5.2. Due to resource constraints, we randomly sample up to 500 cases per task for country-specific indicators, and up to 1000 cases per task for global indicators. In practice, some tasks contain fewer samples due to data availability limitations, but these values represent the maximum sample size allowed per task. The detailed statistics of available data before applying the sampling strategy are presented in Figure 3(b).
@@
Figure 3: (a) 11 indicators in benchmark and their counts. (b) Statistics of dataset.
@@ 2.2 Evaluation Methodologies
We design three distinct paradigms to explore the capabilities of LVLMs in socioeconomic indicator prediction. As shown in Figure 4, each paradigm is aimed at evaluating a different facet of how LVLMs can be applied to this task.
Direct Metric Prediction
Direct Metric Prediction refers to providing region-level urban imagery and directly querying the LVLM for the metric value, such as: “What’s the percentage of the population commuting by public transit in this census tract?”. In addition, the prompt positions the model as an urban socioeconomic scientist in a specific city. Despite this, the model faces significant challenges in accurately predicting the exact true values of these indicators.
Normalized Metric Estimation
Given the difficulty of directly predicting precise indicator values, we adopt a Normalized Metric Estimation approach inspired by GeoLLM (Manvi et al., 2024b). Specifically, we transform all indicator values into a normalized range from 0.0 to 9.9, discretized to one decimal place. The model is then prompted to estimate this normalized value based on the input images. This formulation aims to investigate whether the LVLM possesses coarse-grained spatial knowledge and the ability to associate visual cues with relative indicator levels.
Feature-Based Regression
In the Feature-Based Regression approach, we first design a structured prompt that guides the LVLM to evaluate each street view image along 13 predefined visual attributes, following the visual taxonomy proposed by Fan et al. (2023). These features capture key elements of the urban environment, such as greenery, vehicle, facade, and sidewalk. For each region, we represent its visual environment using 10 sampled street view images. For each visual feature, we compute the average score across these images, resulting in a single feature vector that characterizes the region. These aggregated visual features are then used as inputs to a LASSO regression model, which is trained to predict the ground-truth indicator values using a 5-fold cross-validation setup.
Figure 4: Prompt examples for three evaluation methodologies.
@@ 3 Experiments
3.1 Overall Performance on Feature-Based Regression
Challenge of the Benchmark for LVLMs
As shown in Table 1, the overall performance suggests that our benchmark poses a significant challenge for current large vision-language models. In particular, tasks such as Mental Health and Bachelor Ratio exhibit low $R^{2}$ scores, in some cases even approaching zero, e.g., 0.001. This highlights the difficulty of CityLens in Feature-Based Regression method: even when leveraging visual features extracted by advanced LVLMs, the resulting representations often fail to capture the complex patterns required for accurate prediction of socioeconomic indicators. We also observe that general LVLMs underperform compared to the domain-specific contrastive learning model UrbanVLP on several tasks. Therefore, improving LVLM performance for urban socioeconomic sensing remains an important and open challenge.
@@ 3.1 Overall Performance on Feature-Based Regression
Tasks GDP Pop. HP VC PT DR BH MH AH LE BR Mean SD
[…]
UrbanCLIP 0.450 0.030 0.316 0.033 0.128 0.123 0.612 0.021 0.191 0.024 0.094 0.184 0.196
UrbanVLP 0.717 0.132 0.559 0.149 0.551 0.446 0.807 0.403 0.382 0.025 0.422 0.417 0.243
[…]
@@
Performance Differences Across Models
We observe substantial performance differences across LVLMs, reflecting how model scale, architecture, and training design influence their ability to extract meaningful visual features for downstream prediction. Comparing models within the same series but at different scales, we find that increasing model size does not always guarantee better performance. For example, Gemma3-12B achieves the best score on GDP and Life Expectancy, yet the 27B variant performs worse in these two tasks, with relative drops of 4.3% and 6.8% respectively. This counterintuitive result may be attributed to the unique nature of socioeconomic sensing tasks, which requires the model to consistently extract and score a predefined set of nuanced visual features from urban imagery. When comparing models from different series with similar parameter scales, clear differences emerge. For instance, Gemma3-4B significantly outperforms Qwen2.5VL-3B in nearly all tasks, with relative improvements ranging from 6.4% to 255% across different indicators, suggesting that Gemma’s architecture or training process may enable more consistent and informative scoring of urban visual features, which in turn leads to better performance in socioeconomic prediction.
Variations Across Different Task Types
Performance also varies across task types. Tasks like Building Height, Public Transport, and GDP tend to have relatively higher values across models, with Building Height reaching an $R^{2}$ of 0.590, suggesting that these indicators are associated with more observable visual cues that can be directly captured from street view images. For instance, Building Height is closely linked to the skyline and the vertical structure visible in images; Public Transport usage may be inferred from the presence of bus stops, transit signs, or road markings. In contrast, tasks such as Life Expectancy and Mental Health remain highly challenging, exhibiting low or near-zero predictive scores for many models. These indicators are influenced by latent factors such as lifestyle, stress levels, or social cohesion, which lack clear visual signals in urban imagery. Even if certain proxies exist, such as the presence of graffiti or the amount of green space, they are often subtle or semantically ambiguous, making it hard for LVLMs to interpret reliably and consistently.